### v1

Abstract
Class Incremental Learning (CIL) constitutes a pivotal subfield within continual learning, aimed at enabling models to progressively learn new classification tasks while retaining knowledge obtained from prior tasks. Although previous studies have predominantly focused on backward compatible approaches to mitigate catastrophic forgetting, recent investigations have introduced forward compatible methods to enhance performance on novel tasks and complement existing backward compatible methods. In this study, we introduce an effective-Rank based Feature Richness enhancement (RFR) method, designed for improving forward compatibility. Specifically, this method increases the effective rank of representations during the base session, thereby facilitating the incorporation of more informative features pertinent to unseen novel tasks. Consequently, RFR achieves dual objectives in backward and forward compatibility: minimizing feature extractor modifications and enhancing novel task performance, respectively. To validate the efficacy of our approach, we establish a theoretical connection between effective rank and the Shannon entropy of representations. Subsequently, we conduct comprehensive experiments by integrating RFR into eleven well-known CIL methods. Our results demonstrate the effectiveness of our approach in enhancing novel-task performance while mitigating catastrophic forgetting. Furthermore, our method notably improves the average incremental accuracy across all eleven cases examined.
keywords:
[…]
1 Introduction
Continual learning, the process of continually acquiring and integrating new knowledge, has emerged as a significant challenge in the field of machine learning. In contrast to conventional learning paradigms that focus on static datasets and fixed tasks, continual learning encompasses the dynamic and ever-changing nature of real-world applications. To address this challenge, class incremental learning (CIL) (Rebuffi et al., 2017; Hou et al., 2019; Douillard et al., 2020; Shi et al., 2022), a subfield of continual learning, primarily focuses on developing techniques that enable adaptive learning to accommodate new classes, while minimizing the detrimental impact of knowledge degradation on previously learned classes.
In CIL, training takes place through multiple incremental sessions, each focusing on a distinct subset of classes that do not overlap with the subsets utilized in other sessions. The base session involves training a model from scratch to perform a base task, typically involving a large number of classes. Subsequently, in each novel session, the model is expected to incrementally learn a new novel task, typically involving a smaller number of classes, while retaining its performance on previously learned tasks. The model’s performance is evaluated using the classes of all the previous and current tasks, without access to the task identification information.
The primary objective in CIL is to mitigate the detrimental impact of catastrophic forgetting, which refers to a significant decline in the model’s performance on previous task classes after learning new classes from a subsequent task. To address this challenge, most of the previous works have focused on addressing the forgetting problem in the updated model (i.e., backward compatible approach) (Zhou et al., 2022; Shi et al., 2022). This is commonly achieved by enforcing similarity between the updated model and its predecessor. For instance, weight regularization methods enforce similarity in weights (Kirkpatrick et al., 2017; Zenke et al., 2017; Aljundi et al., 2018; Chaudhry et al., 2018), while knowledge distillation methods enforce similarity in representations (Rebuffi et al., 2017; Li and Hoiem, 2017; Castro et al., 2018; Hou et al., 2019; Wu et al., 2019).
(a) Representation rank
(b) Novel task performance
(c) Catastrophic forgetting
Figure 1: Impact of increasing representation rank during the base session. We conduct an analysis of UCIR models with and without the integration of our method. ResNet-18 model is trained for the CIFAR-100 dataset, utilizing 50 base classes and a split size of 10 classes for each novel session. (a) Effective rank of the feature extractor. (b) Novel task performance in each novel session. (c) The degree of catastrophic forgetting that occurs for the base task.
In contrast, the concept of forward compatible approach, which aims to facilitate the training of the subsequent tasks, has received relatively little attention. Recently, a few studies have introduced methods aimed at learning forward-compatible representations, predominantly by leveraging class information. These methods include employing dual augmentation techniques (class augmentation and semantic augmentation) (Zhu et al., 2021) and enforcing class-wise decorrelations (CwD) (Shi et al., 2022). In few-shot class incremental learning (FSCIL), another method called FACT also implements forward compatible method by assigning virtual prototypes and relying on class information (Zhou et al., 2022).
To enhance the forward compatibility of the representation, this study focuses on representation rank and feature richness. To be precise, we utilize effective rank (Roy and Vetterli, 2007) in lieu of algebraic rank. Effective rank is a continuous-value extension of algebraic rank. In contrast to algebraic rank, it possesses two advantageous properties: differentiability and the ability to effectively manage extremely small singular values. Representation rank is a general property of representations, and it is not constrained to specific types of information such as class information. Therefore, we conjecture that the representation rank can serve as a crucial indicator of the quantity of encoded features in the representation, with higher-rank representations expected to contain rich features that can be beneficial for subsequent tasks. In this work, we substantiate the conjecture with a theorem and empirical investigations. We prove that the Shannon entropy of the representation is maximized when the effective rank is maximized. Because entropy is a quantitative measure of information, larger entropy can be interpreted as richer features. Empirically, we show that effective rank increases as more classes are included in a plain supervised learning and we also show that effective rank increases as unsupervised learning proceeds.
To this end, we propose an effective-Rank based Feature Richness enhancement (RFR) method that increases the effective rank of representation during the base session in order to preserve informative features. Specifically, we highlight the importance of regularizing the feature extractor during the base session, as it offers two main advantages in the process of learning novel sessions. First, the performance of novel tasks can be enhanced by utilizing the rich features encoded by the base task. Second, catastrophic forgetting can be mitigated because novel tasks can leverage the rich features encoded by the base task, leading to minimal modifications to the feature extractor during the learning of novel tasks. In contrast to IL2A (Zhu et al., 2021), which explicitly implements both backward and forward compatible methods across base and novel sessions, RFR achieves two distinct methodological objectives solely through a forward compatible approach, without explicit adoption of backward compatible regularization methods during novel sessions.
We have performed extensive experiments to confirm the effectiveness of RFR for class incremental learning. While we defer the explanation of the full results until Section 4, a glimpse of the experimental results is provided in Figure 1. Our method can effectively increase the representation rank as shown in Figure 1(a), can improve the performance of novel tasks as exhibited in Figure 1(b), and can substantially mitigate catastrophic forgetting as demonstrated in Figure 1(c). These empirical results provide compelling evidence for the effectiveness of our approach in enhancing the feature richness in the representation.
2 Related works
2.1 Backward compatible approaches
Weight regularization methods aim to minimize the weight distance between the feature extractor learned in the previous session and the feature extractor learned in the subsequent session. In this approach, previous works have primarily focused on calculating the importance of each weight to penalize changes of individual weights. To calculate the importance, several methods have been proposed. EWC (Kirkpatrick et al., 2017) proposed a diagonal approximation of the Fisher information matrix. SI (Zenke et al., 2017) and MAS (Aljundi et al., 2018) proposed a path integral approach, which accumulates the changes in weights throughout the entire learning trajectory. RWalk (Chaudhry et al., 2018) combined the Fisher information matrix approach with the path integral approach.
Representation regularization methods aim to prevent forgetting by imposing a penalty on changes in representations. Typically, a regularization is applied during novel sessions, wherein Knowledge Distillation (Hinton et al., 2015) plays a key role. The previous session’s network acts as the teacher, imparting its knowledge to the student network being trained in the novel session. iCaRL (Rebuffi et al., 2017) employs sigmoid output for knowledge distillation, while other methods (Li and Hoiem, 2017; Castro et al., 2018; Wu et al., 2019; Zhao et al., 2020) utilize temperature-scaled softmax outputs. UCIR (Hou et al., 2019) incorporates cosine normalization, less-forget constraint, and inter-class separation to mitigate the adverse effects of the imbalance between previous and new classes. PODNet (Douillard et al., 2020) effectively reduces the difference in pooled intermediate features along the height and width directions through knowledge distillation.
2.2 Forward compatible approaches
In the pursuit of establishing forward compatibility within class incremental learning scenarios, IL2A (Zhu et al., 2021) introduced two distinct augmentation strategies aimed at improving both backward and forward compatibility. The semantic augmentation generates features from stored distribution of previous tasks, and these features are fed to another classification loss for backward compatibility. Simultaneously, the class augmentation method utilizes an augmented class as an auxiliary class to learn transferable and diverse representation, thereby promoting forward compatibility. The Classwise Decorrelation (CwD) (Shi et al., 2022) method was developed to mimic the behavior of an oracle during the base session. Through empirical investigations, it was discerned that resembling the representation distribution patterns akin to those exhibited by the oracle, characterized by a uniform dispersion of eigenvalues across each class, holds the potential to enhance forward compatibility. To realize this, classwise Frobenius norm of representations was strategically employed during the base session, serving as a mechanism to enforce the desired distribution consistency.
If we consider a broader research area beyond class incremental learning, ForwArd Compatible Training (FACT) also introduced a forward compatible approach for few-shot class incremental learning (Zhou et al., 2022). Within the FACT framework, the concept of virtual classes was incorporated during the training of the base session, effectively allocating embedding space to accommodate upcoming classes. FACT integrates pseudo labels and virtual instances to facilitate effective network training generated through the manifold mixup technique. These elements collectively enhanced the network’s adaptability, underscoring FACT’s significance in fostering the seamless integration of new classes while upholding established knowledge.
While these forward-compatible approaches have demonstrated their effectiveness, they are all reliant on class information. Considering that the training data for the base task can encompass considerably many informative features that can be utilized for novel sessions beyond mere class information, regularization of representations in an unsupervised manner could prove advantageous in encoding richer, forward-compatible features.
3 Enhancing feature richness by increasing representation rank
In this section, we present the details of the proposed method. The primary objective of our method is to enhance the feature richness in the representation by increasing the rank of the feature extractor’s output representation during the base session.
3.1 Effective rank
For a set of $N$ samples in a mini-batch, each having an $L_{2}$-normalized representation vector $\bm{h}_{i}\in\mathbb{R}^{d}$ satisfying $||\bm{h}_{i}||_{2}=1$ and $N>d$, the rank of representation matrix $\bm{H}=[\bm{h}_{1},\bm{h}_{2},...,\bm{h}_{N}]^{T}\in\mathbb{R}^{N\times d}$ can be quantified as
$\mathsf{rank}(\bm{H})=\mathsf{rank}(\bm{U}\bm{\Sigma}\bm{V}^{T})=\mathsf{rank}(\bm{\Sigma})=\sum_{i=1}^{d}\bm{1}_{\mathrm{0<\sigma_{i}}},$ (1)
where $\bm{U}\bm{\Sigma}\bm{V}^{T}$ is a singular value decomposition of $\bm{H}$ and $\{\sigma_{i}\}$ are the singular values arranged in a descending order.
The definition of algebraic rank in Eq. (1) exhibits two practical problems. The first problem is that it equally counts all positive singular values regardless of their strength. Therefore, it can be misleading when extremely small $\sigma_{i}$ values exist. For instance, rank is known to be susceptible to noise (Choi et al., 2017). A commonly adopted remedy for this problem is to set a threshold for counting. We call this thresholded rank as $\mathsf{trank}$, and it is defined as
[…]
where $\rho$ is a threshold parameter chosen between 00 and $1$. Eq. (2) quantifies the count of the largest singular values that collectively encompass $\rho$ proportion of the total singular value energy. The $\mathsf{trank}$ provides a straightforward way to avoid the first practical problem, but it is still vulnerable to the second problem. The second problem is the non-differentiable nature of algebraic rank and $\mathsf{trank}$. They are both integer-valued and thus not differentiable. A direct consequence is the difficulty for integrating rank or $\mathsf{trank}$ into an end-to-end learning. An elegant work-around for this problem is known as effective rank (Roy and Vetterli, 2007). Effective rank is defined as
$\mathsf{erank}(\bm{H})\triangleq\exp\left(-\sum_{i=1}^{d}\lambda_{i}\log{\lambda_{i}}\right),$ (3)
where $\lambda_{i}\triangleq\sigma_{i}^{2}/N$ and $\{\lambda_{i}\}$ corresponds to the set of eigenvalues for $\bm{H}^{T}\bm{H}/N$. We note that $\sum_{i=1}^{d}\lambda_{i}=1$ because $\sum_{i=1}^{d}\lambda_{i}=\sum_{i=1}^{d}\sigma_{i}^{2}/N=tr(\bm{H}^{T}\bm{H}/N)=tr(\sum_{i=1}^{N}\bm{h}_{i}\bm{h}_{i}^{T}/N)\\ =\sum_{i=1}^{N}tr(\bm{h}_{i}\bm{h}_{i}^{T})/N=1$. While being continuous, the effective rank is known to satisfy a list of properties (Roy and Vetterli, 2007). In particular, the following lower bound can be derived.
$\mathsf{erank}(\bm{H})\leq\mathsf{rank}(\bm{H}).$ (4)
The logarithm of $\mathsf{erank}$ turns out to be the same as the von Neumann entropy (Nielsen and Chuang, 2002; Wilde, 2013) and its use for learning representation has been extensively studied in (Kim et al., 2023). In our work, we adopt $\mathsf{erank}$ as the main concept and the logarithm of $\mathsf{erank}$ as the main method for controlling feature richness.
3.2 Empirical investigation of rank vs. feature richness
To understand the relation between the three types of rank and feature richness in representation, we have devised two experiments. For the supervised learning experiment, we have controlled the number of classes used for the training and measured the rank values. For the unsupervised learning experiment, we have performed contrastive learning with SimCLR loss (Chen et al., 2020) and measured the rank values as the learning epoch increases. The results are shown in Figure 2. In both experiments, algebraic rank is fixed at the maximum value ($d=512$) because it counts even extremely small singular values. Both $\mathsf{trank}$ and $\mathsf{erank}$, however, increase as more classes are used for plain supervised learning or as training epoch increases for unsupervised learning. For a more comprehensive exploration of the correlation between representation rank and downstream performance in contrastive learning, refer to the analysis provided in (Garrido et al., 2023). As we will show in Section 4, a strong relationship between $\mathsf{erank}$ and feature richness also holds for class incremental learning.
(a) Supervised
(b) Unsupervised
Figure 2: Rank vs. feature richness. ResNet-18 was trained using ImageNet-100 dataset. (a) Supervised learning – as we include more classes in the training that starts from scratch, both $\mathsf{trank}$ and $\mathsf{erank}$ increase. Algebraic rank remains at the maximum value. (b) Unsupervised learning with SimCLR loss – as the unsupervised representation learning proceeds, both $\mathsf{trank}$ and $\mathsf{erank}$ increase. Algebraic rank remains at the maximum value.
3.3 Proposed method
[…]
where $\alpha>0$ is the strength hyper-parameter. The RFR loss is not applied in the novel sessions, and this decision is analyzed in Section 5.2.
3.4 Connection to Shannon entropy
Shannon entropy is a fundamental measure of information and it quantifies the information contained in a random variable (Cover, 1999). Therefore, entropy of representation can serve as a measure of feature richness. Assuming Gaussian distribution, we prove a theoretical connection between the proposed RFR and Shannon entropy.
Theorem 1.
For representation $\bm{h}\in\mathbb{R}^{d}$ that follows a multivariate Gaussian distribution, the entropy of representation is maximized if the effective rank of the representation is maximized.
Proof.
Without loss of generality, we consider normalized representation vectors (i.e., $||\bm{h}||_{2}=1$) that follow a zero-mean multivariate Gaussian distribution (i.e., $\bm{h}\sim\mathcal{N}(0,\Sigma)$), thereby satisfying the condition $tr(\Sigma)=1$.
The proof is based on two parts. In the first part, we prove that the solution for maximizing the representation entropy is $\Sigma=\frac{1}{d}I$. In the second part, we prove that the solution for maximizing effective rank is also $\Sigma=\frac{1}{d}I$.
The entropy of multivariate Gaussian distribution, $\mathcal{N}(\mu,\Sigma)$, can be computed as (Cover, 1999)
$\frac{d}{2}\log{}2\pi e+\frac{1}{2}\log{\det(\Sigma)},$ (7)
where $\det(\Sigma)$ is the determinant of $\Sigma$. By examining the equation, it can be confirmed that entropy is maximized when $\det(\Sigma)$ is maximized under the constraint of $tr(\Sigma)=1$. The solution for this optimization problem is $\frac{1}{d}I$ because of Hadamard’s inequality and inequality of arithmetic and geometric means. First, the following Hadamard’s inequality states that the determinant of a positive definite matrix is less than the product of its diagonal elements.
$\det(\Sigma)\leq\prod_{i}\Sigma_{ii}\text{, with equality iff }\Sigma\text{ is diagonal.}$. (8)
Therefore, all the off-diagonal terms need to be zero to maximize the entropy. Second, $\lambda_{i}$ needs to be equal to $1/d$ for all $i$ – otherwise, the inequality of arithmetic and geometric implies that $\prod_{i}\Sigma_{ii}$ can be increased further while satisfying the sum constraint of $tr(\Sigma)=1$.
The proof for $\Sigma=\frac{1}{d}I$ being the solution for maximizing the effective rank in Eq. (3) is trivial. The logarithm of effective rank is $-\sum_{i=1}^{d}\lambda_{i}\log{\lambda_{i}}$ where $\sum_{i=1}^{d}\lambda_{i}=1$. Because this can be interpreted as the entropy of a probability distribution denoted by $\{\lambda_{i}\}$, it is maximized by the uniform distribution, i.e., $\lambda_{i}=\frac{1}{d}$. ∎
It is noteworthy that the Gaussian assumption on the representation (Kingma and Welling, 2013; Yang et al., 2021) has not only been empirically observed by numerous researchers but also has been theoretically justified in works including (Williams, 1997; Neal, 2012; Lee et al., 2017; Yang, 2019).
4 Experiments
We provide an overview of our experimental settings in Section 4.1. Subsequently, we demonstrate the efficacy of our method for enhancing novel task performance in Section 4.2 and for mitigating catastrophic forgetting in Section 4.3. In Section 4.4, we demonstrate that our method can improve performance of eleven well-known backward-compatible methods. In Section 4.5, further analysis is provided. In Section 4.6, an ablation study is presented. In Section 4.2 and Section 4.3, we focus our analysis on the UCIR method, given its prevalent adoption across recent state-of-the-art methodologies.
Figure 3: Improvements in forward compatibility – overall accuracy at each session is shown for UCIR. Two ResNet-18 models are trained with and without RFR for ImageNet-100 dataset, utilizing 50 base classes under different split sizes (a) 10, (b) 5, and (c) 2 for each novel session. The feature extractors trained by the 50 classes of the base task remain frozen during novel sessions.
Figure 4: Improvements in forward compatibility – novel-task accuracy at each session is shown for UCIR. Results were obtained from the same experiment as in Figure 6.
Figure 5: Improvements in forward compatibility – overall accuracy at each session is shown for UCIR. Two ResNet-18 models are trained with and without RFR for ImageNet-100 dataset, utilizing 50 base classes under different split sizes (a) 10, (b) 5, and (c) 2 for each novel session.
(a) S=10
(b) S=5
(c) S=2
(a) S=10
(b) S=5
(c) S=2
(a) S=10
(b) S=5
(c) S=2
(a) S=10
(b) S=5
(c) S=2
Figure 3: Improvements in forward compatibility – overall accuracy at each session is shown for UCIR. Two ResNet-18 models are trained with and without RFR for ImageNet-100 dataset, utilizing 50 base classes under different split sizes (a) 10, (b) 5, and (c) 2 for each novel session. The feature extractors trained by the 50 classes of the base task remain frozen during novel sessions.
Figure 4: Improvements in forward compatibility – novel-task accuracy at each session is shown for UCIR. Results were obtained from the same experiment as in Figure 6.
Figure 5: Improvements in forward compatibility – overall accuracy at each session is shown for UCIR. Two ResNet-18 models are trained with and without RFR for ImageNet-100 dataset, utilizing 50 base classes under different split sizes (a) 10, (b) 5, and (c) 2 for each novel session.
Figure 6: Improvements in forward compatibility – novel-task accuracy at each session is shown for UCIR. Results were obtained from the same experiment as in Figure 6.
4.1 Settings
Datasets:
We employ CIFAR-100 and ImageNet-100, two widely adopted benchmark datasets in CIL. To maintain consistency with prior studies, we follow the standard class orderings proposed in (Rebuffi et al., 2017) for all evaluations except for the evaluation of PODNet (Douillard et al., 2020) and IL2A (Zhu et al., 2021), for which we utilize the class orderings defined in PODNet and IL2A, respectively. To evaluate the CIL methods, all classes in datasets are divided into multiple tasks. The initial 50 classes are designated for the base task, while the remaining classes are split into novel tasks, where the size of each split is either 10, 5, or 2. In this study, we denote split sizes of 10, 5, and 2 as S=10, S=5, and S=2, respectively.
Implementation Details:
In this study, ResNet-18 (He et al., 2016) is employed as the base model to investigate. To reproduce the results of iCaRL (Rebuffi et al., 2017), LwF (Li and Hoiem, 2017), SI (Zenke et al., 2017), EEIL (Castro et al., 2018), MAS (Aljundi et al., 2018), RWalk (Chaudhry et al., 2018), BiC (Wu et al., 2019), IL2M (Belouadah and Popescu, 2019), and UCIR (Hou et al., 2019), we utilize the open-source codebase provided by FACIL (Masana et al., 2022). For PODNet (Douillard et al., 2020), IL2A (Zhu et al., 2021), and AFC (Kang et al., 2022), we employ their own open-source codebases. To incorporate RFR into the aforementioned methods, we modify their respective loss functions during the training of the base session, as specified by Eq. (6). Meanwhile, the remaining configurations, such as the loss functions applied during the training of novel sessions and the default hyper-parameter settings, remain unchanged. To evaluate the performance of the models, we adopt the standard metric of average incremental accuracy (AIC) proposed in (Rebuffi et al., 2017).
4.2 Superiority in forward compatibility
To investigate forward compatibility improvement, two models are trained for a base task, one with RFR and one without RFR. Then, we compare the novel task performance while keeping the feature extractors frozen and training only the classification heads for the respective novel tasks. The results are shown in Figure 6, and they demonstrate that our method yields improved novel task AIC performance across all three cases, with enhancements of 2.50%, 2.63%, and 2.59% for split sizes of 10, 5, and 2, respectively. To delve deeper into this improvement, we analyze the performance of each individual novel task. As shown in Figure 6, our method leads to enhanced performance for the majority of novel tasks. Moreover, the average performance in novel tasks demonstrates significant enhancements, with improvements of 5.93%, 7.11%, and 7.77% for split sizes of 10, 5, and 2, respectively. These results strongly indicate that our method is effective in promoting forward compatible representations. Additionally, we conduct a similar analysis, this time without freezing of feature extractors. The results presented in Figure 6 and Figure 6 exhibit similar performance enhancements, thereby substantiating superiority of RFR’s forward compatibility.
4.3 Mitigation in catastrophic forgetting
Figure 7: Weight change from the base session for UCIR. Two ResNet-18 models are trained with and without RFR for ImageNet-100 dataset, utilizing 50 base classes under different split sizes (a) 10, (b) 5, and (c) 2 for each novel session.
Figure 8: Weight change from the immediate previous session for UCIR. Two ResNet-18 models are trained with and without RFR for ImageNet-100 dataset, utilizing 50 base classes under different split sizes (a) 10, (b) 5, and (c) 2 for each novel session.
Figure 9: Cosine similarity in representation with respect to the base session for UCIR. Two ResNet-18 models are trained with and without RFR for ImageNet-100 dataset, utilizing 50 base classes under different split sizes (a) 10, (b) 5, and (c) 2 for each novel session.
(a) S=10
(b) S=5
(c) S=2
(a) S=10
(b) S=5
(c) S=2
(a) S=10
(b) S=5
(c) S=2
(a) S=10
(b) S=5
(c) S=2
Figure 7: Weight change from the base session for UCIR. Two ResNet-18 models are trained with and without RFR for ImageNet-100 dataset, utilizing 50 base classes under different split sizes (a) 10, (b) 5, and (c) 2 for each novel session.
Figure 8: Weight change from the immediate previous session for UCIR. Two ResNet-18 models are trained with and without RFR for ImageNet-100 dataset, utilizing 50 base classes under different split sizes (a) 10, (b) 5, and (c) 2 for each novel session.
Figure 9: Cosine similarity in representation with respect to the base session for UCIR. Two ResNet-18 models are trained with and without RFR for ImageNet-100 dataset, utilizing 50 base classes under different split sizes (a) 10, (b) 5, and (c) 2 for each novel session.
Figure 10: Cosine similarity in representation with respect to the immediate previous session for UCIR. Two ResNet-18 models are trained with and without RFR for ImageNet-100 dataset, utilizing 50 base classes under different split sizes (a) 10, (b) 5, and (c) 2 for each novel session.
To analyze improvement in on catastrophic forgetting, we conduct three investigations. First, we examine the influence of our method on the weight changes of the feature extractor. As novel sessions progress, the $L_{2}$-weight distance between the feature extractor learned in the base session and that in the novel session increases, as demonstrated in Figure 10. However, with the integration of RFR, the weight distances are significantly reduced across all sessions. The average reductions in weight distances are 8.70, 8.13, and 6.79 for split sizes of 10, 5, and 2, respectively. Additionally, Figure 10 reveals that the weight distance with the feature extractor from the immediate previous session also increases less when our method is employed, showing average reduction of 6.55, 4.95, and 2.99 for split sizes of 10, 5, and 2, respectively.
Second, we investigate the impact of our method on cosine similarity in representations produced from the validation dataset of the base task. As shown in Figure 10, our method leads to an increase in the similarity between the representations learned in the base session and those in novel sessions. The increase in representation similarity across all sessions is significant, with average increase of 0.09, 0.12, and 0.13 for split sizes of 10, 5, and 2, respectively. Moreover, Figure 10 demonstrates that the similarity with representations from the immediate previous session also increases, with average increase of 0.02, 0.02, and 0.03 for split sizes of 10, 5, and 2, respectively.
(a) S=10
(b) S=5
(c) S=2
Figure 11: Catastrophic forgetting in the base task for UCIR. Two ResNet-18 models are trained with and without RFR for ImageNet-100 dataset, utilizing 50 base classes under different split sizes (a) 10, (b) 5, and (c) 2 for each novel session.
Finally, we evaluate the actual impact on catastrophic forgetting in the base task, measured by the performance drop on the base task (Liu et al., 2020; Mittal et al., 2021). Figure 11 clearly demonstrates that our method leads to decreased catastrophic forgetting in the base task across all sessions, with an average reduction of 1.30%, 2.84%, and 2.90% for split sizes of 10, 5, and 2, respectively.
4.4 Consistent improvements over eleven existing methods
Table 1: Performance improvements by RFR. All methods in this table follow the same class orderings as originally proposed in iCaRL (Rebuffi et al., 2017).
Method CIFAR100 (B=50) ImageNet-100 (B=50)
S=10 S=5 S=2 S=10 S=5 S=2
iCaRL (Rebuffi et al., 2017) 49.83±${}_{\pm 2.33}$ 46.86±${}_{\pm 2.38}$ 44.69±${}_{\pm 1.97}$ 53.77±${}_{\pm 1.68}$ 48.41±${}_{\pm 8.70}$ 49.78±${}_{\pm 3.65}$
  with RFR 50.56±${}_{\pm 1.47}$ 48.48±${}_{\pm 1.17}$ 44.99±${}_{\pm 0.97}$ 57.17±${}_{\pm 1.66}$ 55.62±${}_{\pm 0.71}$ 54.40±${}_{\pm 1.63}$
Improvement +0.72 +1.62 +0.3 +3.4 +7.21 +4.61
LwF (Li and Hoiem, 2017) 39.25±${}_{\pm 2.04}$ 26.53±${}_{\pm 1.87}$ 33.41±${}_{\pm 2.06}$ 54.26±${}_{\pm 1.32}$ 50.30±${}_{\pm 1.96}$ 44.52±${}_{\pm 1.70}$
  with RFR 40.98±${}_{\pm 0.34}$ 28.82±${}_{\pm 4.17}$ 34.46±${}_{\pm 0.63}$ 56.48±${}_{\pm 1.22}$ 52.52±${}_{\pm 0.35}$ 47.57±${}_{\pm 1.02}$
Improvement +1.73 +2.29 +1.06 +2.22 +2.21 +3.05
SI (Zenke et al., 2017) 37.90±${}_{\pm 2.39}$ 22.88±${}_{\pm 1.41}$ 31.01±${}_{\pm 3.24}$ 51.89±${}_{\pm 1.70}$ 49.25±${}_{\pm 2.19}$ 44.16±${}_{\pm 1.64}$
  with RFR 39.41±${}_{\pm 0.60}$ 24.02±${}_{\pm 0.38}$ 33.48±${}_{\pm 0.80}$ 54.99±${}_{\pm 1.08}$ 51.98±${}_{\pm 0.89}$ 47.46±${}_{\pm 1.09}$
Improvement +1.51 +1.13 +2.47 +3.1 +2.73 +3.29
EEIL (Castro et al., 2018) 37.81±${}_{\pm 2.55}$ 24.48±${}_{\pm 1.18}$ 32.37±${}_{\pm 2.52}$ 52.04±${}_{\pm 1.90}$ 48.65±${}_{\pm 2.34}$ 43.95±${}_{\pm 1.65}$
  with RFR 39.52±${}_{\pm 0.38}$ 25.79±${}_{\pm 1.18}$ 33.46±${}_{\pm 0.69}$ 54.74±${}_{\pm 0.91}$ 51.64±${}_{\pm 0.80}$ 47.89±${}_{\pm 0.60}$
Improvement +1.72 +1.31 +1.09 +2.71 +2.98 +3.94
MAS (Aljundi et al., 2018) 38.02±${}_{\pm 2.49}$ 26.45±${}_{\pm 1.18}$ 31.66±${}_{\pm 3.20}$ 52.17±${}_{\pm 1.75}$ 50.01±${}_{\pm 2.27}$ 46.61±${}_{\pm 1.15}$
  with RFR 39.54±${}_{\pm 0.71}$ 28.80±${}_{\pm 4.52}$ 33.36±${}_{\pm 0.52}$ 55.04±${}_{\pm 1.30}$ 52.77±${}_{\pm 1.02}$ 50.32±${}_{\pm 1.38}$
Improvement +1.53 +2.35 +1.71 +2.88 +2.76 +3.71
RWalk (Chaudhry et al., 2018) 35.75±${}_{\pm 1.63}$ 23.34±${}_{\pm 2.06}$ 27.02±${}_{\pm 5.36}$ 41.08±${}_{\pm 1.91}$ 21.44±${}_{\pm 5.69}$ 20.81±${}_{\pm 2.12}$
  with RFR 39.00±${}_{\pm 1.15}$ 24.44±${}_{\pm 1.06}$ 32.85±${}_{\pm 1.40}$ 44.83±${}_{\pm 1.87}$ 35.42±${}_{\pm 1.71}$ 22.68±${}_{\pm 4.70}$
Improvement +3.25 +1.1 +5.82 +3.75 +13.98 +1.88
BiC (Wu et al., 2019) 54.36±${}_{\pm 1.23}$ 43.42±${}_{\pm 2.70}$ 26.73±${}_{\pm 1.89}$ 61.00±${}_{\pm 4.69}$ 53.12±${}_{\pm 5.68}$ 25.44±${}_{\pm 1.74}$
  with RFR 58.06±${}_{\pm 0.73}$ 48.02±${}_{\pm 2.33}$ 27.74±${}_{\pm 3.11}$ 64.18±${}_{\pm 1.94}$ 53.77±${}_{\pm 3.05}$ 27.56±${}_{\pm 2.72}$
Improvement +3.71 +4.60 +1.01 +3.18 +0.65 +2.13
IL2M (Belouadah and Popescu, 2019) 40.85±${}_{\pm 2.22}$ 24.92±${}_{\pm 1.12}$ 33.41±${}_{\pm 2.12}$ 56.78±${}_{\pm 1.34}$ 52.35±${}_{\pm 2.16}$ 47.10±${}_{\pm 1.85}$
  with RFR 42.54±${}_{\pm 0.49}$ 25.19±${}_{\pm 0.41}$ 35.20±${}_{\pm 0.78}$ 58.40±${}_{\pm 1.04}$ 55.25±${}_{\pm 0.78}$ 50.09±${}_{\pm 0.83}$
Improvement +1.69 +0.27 +1.78 +1.62 +2.9 +2.99
UCIR (Hou et al., 2019) 66.30±${}_{\pm 0.36}$ ${}_{\pm 0.56}$ 52.74±${}_{\pm 0.72}$ 70.57±${}_{\pm 0.51}$ 67.62±${}_{\pm 0.37}$ 63.22±${}_{\pm 0.42}$
  with RFR 69.45±${}_{\pm 0.29}$ 66.16±${}_{\pm 0.10}$ 61.23±${}_{\pm 0.13}$ 71.65±${}_{\pm 0.52}$ 69.52±${}_{\pm 0.13}$ 65.46±${}_{\pm 0.57}$
Improvement +3.14 +5.59 +8.49 +1.08 +1.90 +2.24
AFC (Kang et al., 2022) 70.35±${}_{\pm 0.29}$ 67.02±${}_{\pm 0.24}$ 65.05±${}_{\pm 0.27}$ 77.10±${}_{\pm 0.03}$ 75.27±${}_{\pm 0.18}$ 72.90±${}_{\pm 0.44}$
  with RFR 70.85±${}_{\pm 0.21}$ 67.82±${}_{\pm 0.37}$ 65.53±${}_{\pm 0.24}$ 77.19±${}_{\pm 0.08}$ 75.47±${}_{\pm 0.23}$ 72.99±${}_{\pm 0.04}$
Improvement +0.50 +0.80 +0.48 +0.09 +0.20 +0.09
Average Improvement +1.95 +2.11 +2.42 +2.40 +3.75 +2.79
(a) S=10
(b) S=5
(c) S=2
Figure 12: Overall accuracy at each session for UCIR. Two ResNet-18 models are trained with and without RFR for CIFAR-100 dataset, utilizing 50 base classes under different split sizes (a) 10, (b) 5, and (c) 2 for each novel session.
Table 2: Performance improvements by RFR for PODNet. Unlike the other works, PODNet used its own class orderings for evaluation. As an effort to make comparisons as fair as possible, we followed PODNet’s class orderings for this table.
Method CIFAR100 (B=50) ImageNet-100 (B=50)
S=10 S=5 S=2 S=10 S=5 S=2
PODNet(NME) (Douillard et al., 2020) 68.47±${}_{\pm 1.27}$ 67.09±${}_{\pm 1.19}$ 65.16±${}_{\pm 1.02}$ 60.42±${}_{\pm 0.53}$ 50.93±${}_{\pm 0.73}$ 36.36±${}_{\pm 0.30}$
  with RFR 69.51±${}_{\pm 0.56}$ 67.66±${}_{\pm 0.63}$ 65.59±${}_{\pm 0.48}$ 64.08±${}_{\pm 0.45}$ 56.34±${}_{\pm 0.75}$ 41.03±${}_{\pm 2.15}$
Improvement +1.04 +0.57 +0.43 +3.66 +5.42 +4.67
PODNet(CNN) 66.11±${}_{\pm 0.63}$ 63.49±${}_{\pm 0.45}$ 59.61±${}_{\pm 0.53}$ 73.76±${}_{\pm 0.17}$ 68.32±${}_{\pm 0.27}$ 61.82±${}_{\pm 0.68}$
  with RFR 67.13±${}_{\pm 0.86}$ 64.72±${}_{\pm 1.01}$ 61.49±${}_{\pm 0.71}$ 74.42±${}_{\pm 0.27}$ 69.66±${}_{\pm 0.09}$ 63.11±${}_{\pm 0.22}$
Improvement +1.02 +1.23 +1.89 +0.66 +1.34 +1.29
Average Improvement +1.03 +0.90 +1.16 +2.16 +3.38 +2.98
To demonstrate the efficacy of RFR in enhancing performance, we conduct comprehensive experiments over eleven well-known existing works. The results presented in Table 1 and Table 2 provide compelling evidence of the significant and consistent performance improvements achieved through the integration of RFR into the previous works. Specifically, we observe notable average performance improvements for the state-of-the-art works, including a 3.74% average improvement for UCIR, a 2.63% average improvement for PODNet (NME), a 1.24% average improvement for PODNet (CNN), and a 0.36% average improvement for AFC. Particularly, a remarkable performance improvement is observed in the case of UCIR for CIFAR-100 and its further analysis is shown in Figure 12. The achieved improvements are 3.14% increase for S=10, 5.59% for S=5, and 8.49% for S=2.
Table 3: Performance improvements by RFR for a non-exemplar based Class-IL method, IL2A (Zhu et al., 2021). All results are trained using CIFAR-100 with its own class orderings.
Method S=10 S=5 S=2
IL2A 65.53±${}_{\pm 0.46}$ 58.58±${}_{\pm 0.76}$ 54.31±${}_{\pm 0.54}$
  with RFR 66.07±${}_{\pm 0.33}$ 60.06±${}_{\pm 0.48}$ 55.76±${}_{\pm 0.12}$
Improvement +0.54 +1.48 +1.45
Table 4: Performance improvements by RFR with larger-scale dataset, ImageNet-1000.
Method ImageNet-1000 (B=500)
S=100 S=50
UCIR 65.38±${}_{\pm 0.04}$ 63.44±${}_{\pm 0.11}$
  with RFR 65.69±${}_{\pm 0.10}$ 63.64±${}_{\pm 0.25}$
Improvement +0.31 +0.20
Table 5: Performance improvements by RFR with larger-scale models, ResNet-50 and ViT-Small.
Model Method S=10 S=5 S=2
ResNet-50 UCIR 67.10±${}_{\pm 0.20}$ 61.42±${}_{\pm 0.61}$ 51.62±${}_{\pm 0.14}$
  with RFR 68.57±${}_{\pm 0.12}$ 64.81±${}_{\pm 0.15}$ 57.71±${}_{\pm 0.31}$
Improvement +1.47 +3.39 +6.09
ViT-Small UCIR 50.45±${}_{\pm 0.35}$ 49.03±${}_{\pm 0.27}$ 46.24±${}_{\pm 0.79}$
  with RFR 51.86±${}_{\pm 0.35}$ 50.50±${}_{\pm 0.47}$ 47.96±${}_{\pm 1.03}$
Improvement +1.41 +1.47 +1.72
4.5 Further analysis
Non-exemplar-based approach:
While the effectiveness of RFR is demonstrated for the most popular CIL methods in Section 4.4, they all belong to exemplar-based approaches. Therefore, we further conduct an analysis of RFR for the non-exemplar-based CIL approach, IL2A (Zhu et al., 2021). The results presented in Table 3 exhibit significant performance enhancements, indicating an average improvement of 1.16%.
Larger-scale dataset and models:
In addition to evaluating RFR’s performance under the widely used benchmark settings, we extend our evaluation to a larger-scale dataset. As shown in Table 4, RFR continues to demonstrate its effectiveness for ImageNet-1000 dataset. Furthermore, we analyze RFR’s performance with larger-scale models. The results presented in Table 5 underscore RFR’s efficacy across a range of larger-scale models, including ResNet-50 and ViT-Small (Lee et al., 2021).
4.6 Ablation study
Sensitivity study of regularization coefficient:
We performed a sensitivity study of the strength hyper-parameter $\alpha$ in Eq. (6). The results are presented in Figure 13, and they demonstrate consistent and smooth inverted U-shaped patterns in performance, with the peak performance observed at the value of $\alpha=0.1$.
(a) S=10
(b) S=5
(c) S=2
Figure 13: Impact of regularization coefficient on the performance. ResNet-18 models are trained with a range of regularization coefficients for ImageNet-100 dataset, utilizing 50 base classes under different split sizes (a) 10, (b) 5, and (c) 2 for each novel session.
Impact of the number of classes in the base task:
We demonstrate the consistent effectiveness of RFR across various numbers of classes in the base task. The results are presented in Table 6. Specifically, RFR’s effectiveness is relatively lower when B=10, potentially attributable to the limited features within the 10 classes. However, as the number of classes in the base task increases, the effectiveness of RFR also increases, likely due to the expanded availability of features within the larger number of classes.
Table 6: Performance improvements by RFR for various number of classes in the base task. All results are trained using CIFAR-100 with the same class orderings as originally proposed in iCaRL (Rebuffi et al., 2017).
Classes in base tasks UCIR with RFR Improvement
B=10 60.31±${}_{\pm 0.22}$ 60.41±${}_{\pm 0.34}$ +0.10
B=20 63.16±${}_{\pm 0.34}$ 65.11±${}_{\pm 0.29}$ +1.95
B=30 65.17±${}_{\pm 0.25}$ 66.98±${}_{\pm 0.21}$ +1.81
B=40 67.22±${}_{\pm 0.25}$ 69.53±${}_{\pm 0.12}$ +2.31
B=50 66.30±${}_{\pm 0.36}$ 69.45±${}_{\pm 0.29}$ +3.15
B=60 70.44±${}_{\pm 0.13}$ 73.06±${}_{\pm 0.36}$ +2.62
B=70 71.57±${}_{\pm 0.39}$ 74.20±${}_{\pm 0.20}$ +2.63
B=80 73.07±${}_{\pm 0.18}$ 75.95±${}_{\pm 0.09}$ +2.88
B=90 73.41±${}_{\pm 0.43}$ 75.98±${}_{\pm 0.19}$ +2.57
5 Discussion
5.1 Comparison with CwD
Recently, CwD (Shi et al., 2022), another forward-compatible method, has improved previous state-of-the-art methods in CIL by enforcing class-wise decorrelation. Because both RFR and CwD regularize representation during the base session, a comparative analysis of the two methods was conducted. Specifically, CwD regularizes the Frobenius norm of representation within the same class, while RFR regularizes the eigenvalues of representation without considering class information. However, it has been demonstrated that regularizing the eigenvalues of representation is more effective than regularizing the Frobenius norm in (Kim et al., 2023). The effectiveness of RFR is further confirmed through subsequent evaluation. The performance evaluation was carried out using the UCIR and PODNet models with the CIFAR-100 dataset. As shown in Table 7, RFR consistently demonstrates superior performance, with significant improvements of 1.76%, 2.27%, and 3.10% on average for split sizes of 10, 5, and 2, respectively. These improvements tend to be larger when dealing with smaller split sizes, or equivalently, when handling a larger number of novel sessions.
Table 7: Comparison with CwD.
CIFAR100 (B=50)
UCIR PODNet (CNN)
S=10 S=5 S=2 S=10 S=5 S=2
Baseline 66.30±${}_{\pm 0.36}$ ${}_{\pm 0.56}$ 52.74±${}_{\pm 0.72}$ 66.11±${}_{\pm 0.63}$ 63.49±${}_{\pm 0.45}$ 59.61±${}_{\pm 0.53}$
  with CwD 67.06±${}_{\pm 0.12}$ 62.71±${}_{\pm 0.36}$ 56.39±${}_{\pm 0.30}$ 66.01±${}_{\pm 0.75}$ 63.63±${}_{\pm 1.02}$ 60.14±${}_{\pm 1.16}$
  with RFR 69.45±${}_{\pm 0.29}$ 66.16±${}_{\pm 0.10}$ 61.23±${}_{\pm 0.13}$ 67.13±${}_{\pm 0.86}$ 64.72±${}_{\pm 1.01}$ 61.49±${}_{\pm 0.71}$
Difference +2.39 +3.45 +4.84 +1.12 +1.09 +1.35
5.2 Increasing representation rank during novel sessions
RFR increases effective rank during the base session only. It is also possible to increase effective rank during novel sessions with the goal of acquiring additional features that may be useful in the subsequent novel sessions. Such a strategy, however, can also intensify the modifications of the learned model from the previous sessions. To investigate the overall effect, we have conducted an experiment and the results are shown in Table 8. While RFR demonstrates effectiveness in enhancing feature richness during the base session, its influence on overall performance is marginal when also applied to novel sessions. Therefore, we have chosen to apply effective rank regularization only during the base session.
Table 8: Influence of increasing representation rank during novel sessions.
CIFAR100 (B=50)
S=10 S=5 S=2
Base Only Base+Novel Diff. Base Only Base+Novel Diff. Base Only Base+Novel Diff.
BiC 58.06±${}_{\pm 0.73}$ 58.29±${}_{\pm 0.38}$ +0.23 48.02±${}_{\pm 2.33}$ 46.60±${}_{\pm 1.91}$1.41 27.74±${}_{\pm 3.11}$ 27.64±${}_{\pm 1.83}$0.10
EEIL 39.52±${}_{\pm 0.38}$ 39.44±${}_{\pm 0.63}$0.08 25.79±${}_{\pm 1.18}$ 25.62±${}_{\pm 1.10}$0.17 33.46±${}_{\pm 0.69}$ 33.16±${}_{\pm 1.07}$0.29
iCaRL 50.56±${}_{\pm 1.47}$ 50.50±${}_{\pm 2.04}$0.06 48.48±${}_{\pm 1.17}$ 48.23±${}_{\pm 1.14}$0.25 44.99±${}_{\pm 0.97}$ 44.81±${}_{\pm 1.18}$0.18
IL2M 42.54±${}_{\pm 0.49}$ 42.76±${}_{\pm 0.83}$ +0.21 25.19±${}_{\pm 0.41}$ 26.22±${}_{\pm 1.77}$ +1.03 35.20±${}_{\pm 0.78}$ 34.87±${}_{\pm 0.67}$0.32
LwF 40.98±${}_{\pm 0.34}$ 40.69±${}_{\pm 0.81}$0.28 28.82±${}_{\pm 4.17}$ 26.70±${}_{\pm 0.39}$2.12 34.46±${}_{\pm 0.63}$ 34.63±${}_{\pm 0.75}$ +0.17
MAS 39.54±${}_{\pm 0.71}$ 39.77±${}_{\pm 0.66}$ +0.23 28.80±${}_{\pm 4.52}$ 29.03±${}_{\pm 3.50}$ +0.23 33.36±${}_{\pm 0.52}$ 33.91±${}_{\pm 0.91}$ +0.55
SI 39.41±${}_{\pm 0.60}$ 39.51±${}_{\pm 0.58}$ +0.10 24.02±${}_{\pm 0.38}$ 23.89±${}_{\pm 0.19}$0.13 33.48±${}_{\pm 0.80}$ 33.60±${}_{\pm 0.52}$ +0.12
RWalk 39.00±${}_{\pm 1.15}$ 38.91±${}_{\pm 1.39}$0.09 24.44±${}_{\pm 1.06}$ 25.96±${}_{\pm 3.06}$ +1.53 32.85±${}_{\pm 1.40}$ 33.04±${}_{\pm 1.20}$ +0.19
UCIR 69.45±${}_{\pm 0.29}$ 69.57±${}_{\pm 0.23}$ +0.12 66.16±${}_{\pm 0.10}$ 66.55±${}_{\pm 0.16}$ +0.39 61.23±${}_{\pm 0.13}$ 61.27±${}_{\pm 0.29}$ +0.04
Average difference +0.04 -0.10 +0.02
6 Conclusion
In this study, we propose an effective-Rank based Feature Richness (RFR) enhancement method that can be integrated with a wide range of existing methods in CIL. More specifically, our method increases the effective rank of representations during the base session. In order to substantiate the effectiveness of our method, we have established a theoretical connection between the effective rank and the Shannon entropy of representations. Through empirical analysis, we have demonstrated the efficacy of our method in several dimensions including enhancement in forward compatibility, mitigation of catastrophic forgetting, and improvement in performance. In summary, our method effectively enhances the performance by increasing effective rank and feature richness in the representation.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have influenced the work reported in this paper.
Acknowledgement
This work was supported by a National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. NRF-2020R1A2C2007139) and in part by Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [NO.2021-0-01343, Artificial Intelligence Graduate School Program (Seoul National University)] and Basic Science Research Program through the National Research Foundation of Korea(NRF) funded by the Ministry of Education(NRF-2022R1A6A1A03063039).